OpenStudy (anonymous):

you swing a 4.7-kg bucket of water in a vertical cicle of the radius 3.2m. What spped must the bucket have if it is to complete the circle

OpenStudy (souvik):

the bucket must have a speed \[v \ge \sqrt{gr}\] g = gravitational acceleration r= the radius of the circle here g=9.8 m/s^2 and r=3.2m \[v \ge 5.6 m/s\]